Q about multiple dilutions - I have 9L of old motor oil; I replace 7L with new oil and mix. Then I replace 4.5L with new oil again. What is the final dilution of old oil remaining in the final 9L?


1
Expert's answer
2022-03-16T10:18:04-0400

After the first replacement: 9 – 7 = 2 L of old oil left.

2/7 = 0.2857 or 28.57% of all the volume.

When 4.5 L of this mixture was replaced again, 4.5 L left. So, it’s a double dilution:

28.57 / 2 = 14.3%

So, only 14.3% remains.